# Bidimensional Exploration of the warm-Temperature Ionised gaS (BETIS)

### v1

@@
The extraplanar diffuse ionised gas (eDIG) is a key component for understanding the feedback processes that connect galactic discs and their halos. In this paper, we present the second study of the Bidimensional Exploration of the warm-Temperature Ionised Gas (BETIS) project, which aims to explore the possible ionisation mechanisms and characteristics of the eDIG. We use a sample of eight edge-on galaxies observed with the Multi-Unit Spectroscopic Explorer (MUSE) integral field spectrograph (IFS) and apply the methodology for obtaining binned emission line maps developed in the first paper of the BETIS project. We found that the vertical and radial profiles of the [N II] /H , [S II] /H , [O III] /H , and [O I] /H ratios depict a complex ionisation structure within galactic halos, influenced by the spatial distribution of H II regions across the galactic plane as observed from our line of sigh, with Lyman continuum photon leakage from OB associations constituting the main ionisation source. Moreover, the electron temperature and S+/S ionisation ratio also exhibit dependency on the distribution of H II regions within the galactic discs. Our analysis excludes low-mass, hot, and evolved stars (HOLMES) as viable candidates for secondary ionisation sources to elucidate the unusual behaviour of the line ratios at greater distances from the galactic midplane. In contrast, we ascertain that shocks induced in the interstellar medium by star formation (SF) related feedback mechanisms represent a promising secondary ionisation source of the eDIG. We present a suite of models integrating ionisation mechanisms arising from fast shocks and photoionisation associated with star formation. When applied to the classical Baldwin-Phillips-Terlevich (BPT) diagrams, these models reveal that the ionisation budget of the eDIG ranges from 20% to 50% across our sample, with local variations of up to 20% within individual galaxy halos. This contribution correlates with the presence of filaments and other structural components observed within galaxy halos. The presence of shocks is additionally supported by the observation of high-density, high [O I] /H ratios, characteristic of shock-compressed ionised gas, likely induced by feedback from regions of intense SF within the galactic disk. These results demonstrate consistency across all galaxies analysed in this sample.
@@
To understand the physical mechanisms involved in feedback processes and the baryon cycle, which are fundamental to galaxy evolution it is essential to study the different gas phases present in the interstellar medium (ISM, Ceverino & Klypin, 2008, 2009; Hopkins et al., 2014; Klessen & Glover, 2016; Gatto et al., 2017; Grisdale, 2017; Barnes et al., 2023). One of these gas phases is known as warm ionised medium (WIM, in the context of the Milky Way) or diffuse ionised gas (DIG, for extragalactic sources; Hoyle & Ellis 1963; Hewish et al. 1968; Reynolds 1971; Reynolds et al. 1973; Guélin 1974; Reynolds 1985, 1989; Kennicutt et al. 1989; Ferguson et al. 1996). This warm ( K) and low density (cm-3) gas represents one of the most prominent components of the ISM, constituting the 20% of the ISM in volume and the 90% of the ionising hydrogen in mass (Greenawalt et al., 1997; Wang et al., 1997; Zurita et al., 2000; Otte et al., 2001; Finkbeiner, 2003; Hoopes & Walterbos, 2003; Haffner et al., 2009a). In particular, the DIG component in galactic halos, also known as extraplanar diffuse ionised gas (eDIG), plays a crucial role in the physical processes driving the exchange of gas, metals, and energy between the galactic disk and halo, influencing the evolution of disk galaxies (Tenorio-Tagle & Bodenheimer, 1988; Rand et al., 1990; Dettmar, 1990; Miller & Veilleux, 2003). Studies of the eDIG through imaging encompass explorations using different wavelengths (Sardaneta et al., 2024). In the optical range, the first detections of a diffuse component in the halo were made using H images (Dettmar, 1990; Rand et al., 1990; Pildis et al., 1994; Rand, 1997, 2000; Tacchella et al., 2022), These studies found that in a wide sample of star-forming galaxies the eDIG H emission in the halo correlates with the star formation rate (Rossa & Dettmar, 2003a, b; Jo et al., 2018; Lu et al., 2023). Additionally, sub-arcsecond resolution H imaging studies of individual galaxies (Howk & Savage, 2000; Rossa et al., 2004, 2005) reveal that the eDIG exhibit complex morphological structures such as filaments, superbubbles or chimneys (also detected in X-rays and radio-emission; Tüllmann et al. 2000, 2006). Alternatively, long-slit spectroscopy studies (in combination with photometric data) have enabled kinematic analyses and investigations into the potential ionisation mechanisms of the eDIG. The behaviour of the main ionisation species reveals that the [N II] /H , [S II] /H , [O III] /H and [O I] /H ratios tends to increase as a function of the distance from the midplane (Rand et al., 1990; Dettmar, 1998; Rand, 1998, 2000; Collins & Rand, 2001). Additionally, this higher degree of ionisation of the collisional lines usually implies higher electron temperatures in the eDIG compared to the H II regions (Haffner et al., 1999, 2009b; Boettcher et al., 2019). To explain these features and the increase in the emission of high ionisation species, several additional heating sources have been proposed (Haffner+ 2009 and references therein). To infer the possible physical processes and the roles played by different morphological structures in the halo leading to the eDIG, the simultaneous spatial and spectral information provided by integral field spectrographs (IFS) has resulted in significant progress in this area. One of these studies, carried by the Sydney-AAO Multi-object Integral field spectrograph (SAMI; Croom et al. 2012; Bryant et al. 2015), used a sample of 40 low resolution edge-on galaxies (0.8 kpc ¡ 2.8 kpc) to perform a systematical exploration of the eDIG (Ho et al., 2016) noticing that the disc-halo interaction can influence the eDIG emission through starburst-driven winds. On the other hand, studies with the Mapping Nearby Galaxies at Apache Point (MaNGA; Bundy et al. 2015) IFS using also low resolution large samples found an increment of the [N II] /H indicating a temperature gradient in the halo, as well as a flattening in the H surface brightness ( ) comparable with the [O II] emission at higher distances (6 kpc).
@@
The most recent studies with the Calar Alto Legacy Integral Field Area Survey (CALIFA; Sánchez et al. 2012, 2016) studied the eDIG kinematically using a sample of 25 galaxies with spatial resolution of 0.8 kpc on average (Levy et al., 2019), finding a ”lag” in the rotation velocity of the ionised gas, i.e., a decrement of the radial velocity with the increasing of the distance from the midplane (also reported by other authors, e.g., Rand 2000; Miller & Veilleux 2003; Bizyaev et al. 2017). Nevertheless, they did not find correlations between the radial variation of the lag and the origin of the extraplanar gas. However, performing diagnoses of the main line ratios, they identified the primary ionisation sources of the eDIG the leaking of Lyman continuum (Lyc) photons from the OB stars of the H II regions, with a low contribution of an additional source in the form of hot low mass evolved stars (HOLMES; Flores-Fajardo et al. 2011; Rautio et al. 2022.
@@
In this paper we present the second study of the Bidimensional Exploration of the warm-Temperature Ionised gaS (BETIS) project, focused in the characterisation of the eDIG and the exploration of its possible ionisation mechanisms for a sample of edge-on galaxies. The paper is structured as follows. Section 2 presents and characterises the sample of edge-on galaxies, as well as describes the procedure followed for adaptive binning and the extraction of emission line features, based on the methodology developed in the first part of the series (González-Díaz et al. 2023, BETIS I henceforth). Section 3 shows a description of the ionisation structure and features found in the halo of IC1553 and the connection of the eDIG with the disc for all the sample. Section 4 presents a set of star formation and fast shocks hybrid models to explain the ionisation structure found the halo. In section 5 we show the results of applying the methodology and models of section 4 for the rest of the sample and finally, section 6 outlines the main findings in this paper.
@@
Galaxy | RA (J2000) | DEC (J2000) | Type | z | D | P.A. | FWHM | Incl. | log SB lim. | Bin size |
(deg) | (deg) | (Mpc) | (°) | (pc) | (°) | (erg/s/kpc2) | (pc) | |||
IC217 | 34.0435 | -11.9267 | Scd | 0.00630 | 24.32 | 35.7 | 117.91 | 82.6 | 35.6 | 94.3 |
PGC28308 | 147.5582 | -12.0576 | Scd | 0.00907 | 45.22 | 125.2 | 153.46 | 85.5 | 35.2 | 175.4 |
PGC30591 | 156.3603 | -15.3492 | Sd | 0.00676 | 35.22 | 169.2 | 187.83 | 86.6 | 36.0 | 136.6 |
ESO544-27 | 33.2277 | -19.3168 | Sb | 0.00818 | 32.81 | 153.3 | 159.07 | 90.0 | 35.0 | 127.2 |
IC1553 | 8.1671 | -25.6075 | Irr | 0.00979 | 32.25 | 15.0 | 171.99 | 78.6 | 36.5 | 125.1 |
ESO443-21 | 194.9412 | -29.6002 | Scd | 0.00941 | 40.31 | 160.8 | 195.45 | 79.0 | 36.6 | 156.4 |
ESO469-15 | 347.2317 | -30.8579 | Sb | 0.00545 | 19.81 | 149.2 | 86.44 | 90.0 | 35.9 | 76.9 |
ESO157-49 | 69.9036 | -53.0126 | Sc | 0.00559 | 24.82 | 30.4 | 96.26 | 79.3 | 35.3 | 96.3 |
![Refer to caption](extracted/5688766/IC217_eDIG_map.jpg)
![Refer to caption](extracted/5688766/PGC28308_eDIG_map.jpg)
![Refer to caption](extracted/5688766/PGC30591_eDIG_map.jpg)
![Refer to caption](extracted/5688766/IC1553_eDIG_map.jpg)
![Refer to caption](extracted/5688766/ESO157-49_eDIG_map.jpg)
![Refer to caption](extracted/5688766/ESO443-21_eDIG_map.jpg)
![Refer to caption](extracted/5688766/ESO469-15_eDIG_map.jpg)
![Refer to caption](extracted/5688766/ESO544-27_eDIG_map.jpg)
@@ 2 The data
For this reason, as in BETIS I, we use MUSE IFS data to undertake this task. In particular, the sample of Com19 aligns well with the previous criteria. The sample, listed in table 1, consists on eight edge-on (i ¿ 75°) nearby galaxies (z ¡ 0.01, D ¡ 46 Mpc) with high spatial resolution (FWHM ¡ 200 pc) from the 096.B-0054(A) and 097.B-0041(A) VLT programmes. Table 2 shows the integrated star formation rate (SFR) in the disc spans between 0.024 M⊙/yr ¡ SFR ¡ 0.229 M⊙/yr. It was obtained using the Kennicutt & Evans (2012) relation: log(SFR) = log() - 41.27, with as the H luminosity, corrected for interstellar extinction assuming the Cardelli extinction law, assuming RV = 3.1 (Cardelli et al., 1989) and a Balmer decrement H /H = 2.87. The galaxies are then normal-disc galaxies with low star formation. This is consistent with the classification of these galaxies in the star formation main sequence, with specific SFR ranging between -10.5 ¡ log(sSFR/yr)¡ -9.92, corresponding barely to star-forming galaxies (using MUSE data and 3.6 m stellar masses; Muñoz-Mateos et al. 2015; Zaragoza-Cardiel et al. 2020; Rautio et al. 2022).
@@
This dataset has previously been used to analyse the eDIG in prior studies. Rautio et al. (2022) employed five of these galaxies, in conjunction with H and GALEX-FUV imaging, while Dirks et al. (2023) exclusively utilised IC1553 for a more comprehensive examination of the eDIG ionisation mechanisms within this galaxy. In our paper, we use all eight galaxies from this dataset111We downloaded the reduced datacubes directly from the ESO archive science portal. (eBETIS sample henceforth) to perform an exhaustive spectral and spatial exploration of the eDIG structure and ionisation for every galaxy in particular.
To conduct this exploration, we first extract the emission line features of the eBETIS sample in the form of binned emission line maps, following the same procedure described in BETIS I: we employ our own adaptive binning algorithm to the original datacubes, setting a S/N target of 10 based on [S II] , since this line is one of its main features of the eDIG and DIG in general, as well as it typically exhibits low S/N in both planar DIG and eDIG. Subsequently, we perform a simple stellar population (SSP) synthesis to the binned datacube using STARLIGHT (Cid Fernandes et al., 2005) to the integrated spectra of each bin. This is followed by Gaussian fitting to every emission line of interest present in each resulting nebular spectrum (see BETIS I for more details). This procedure finally results in the binned emission line maps for the H , H , [O III] , [N II] , [S II] and [O I] lines. The average sensitivity of the sample is 35.7 log(erg/s/kpc2), obtained as the 2 of the distribution of H surface brightness ( ) of the bins of every H map. The average physical size of the bins is below 175 pc. Since the average PSF FWHM of the sample is below 200 pc, the binning does not significantly restrict the resolution in our galaxies.
Figure 1 shows the binned H maps of the eBETIS sample. For aesthetic purposes and to highlight the contrast between the galactic plane and the complex structure of the extraplanar gas, the plane of the galaxies is shown in greyscale. To define the galactic plane, we construct an elliptical mask by conducting a 2D Gaussian fitting to the binned H image using the python routine scypy.opt.curve_fit 222scypy.opt.curve_fit documentation., and centring the ellipse in the coordinates of the centre of the galaxy, which also coincides with the centre of the rotation curve of the galaxy. The midplane is then set at z = 0, that corresponds to the centre of the minor axis of the ellipse fitted. The integrated H luminosity of the eDIG for all the galaxies is listed in Table 2. It was obtained integrating the H flux outside the ellipses previously described, and we obtain an average value of 2.51039 erg/s, that is consistent with values obtained by previous authors (Rand, 2000; Miller & Veilleux, 2003; Jo et al., 2018).

### v2

@@
The extraplanar diffuse ionised gas (eDIG) is a key component for understanding the feedback processes that connect galactic discs and their halos. In this paper, we present the second study of the Bidimensional Exploration of the warm-Temperature Ionised Gas (BETIS) project, which aims to explore the possible ionisation mechanisms and characteristics of the eDIG. We use a sample of eight edge-on galaxies observed with the Multi-Unit Spectroscopic Explorer (MUSE) integral field spectrograph (IFS) and apply the methodology for obtaining binned emission line maps developed in the first paper of the BETIS project. We found that the vertical and radial profiles of the [N II] /H , [S II] /H , [O III] /H , and [O I] /H ratios depict a complex ionisation structure within galactic halos, influenced by the spatial distribution of H II regions across the galactic plane as observed from our line of sight, with Lyman continuum photon leakage from OB associations constituting the main ionisation source. Moreover, the electron temperature and S+/S ionisation ratio also exhibit dependency on the distribution of H II regions within the galactic discs. Our analysis excludes low-mass, hot, and evolved stars (HOLMES) as viable candidates for secondary ionisation sources to elucidate the unusual behaviour of the line ratios at greater distances from the galactic midplane. In contrast, we ascertain that shocks induced in the interstellar medium by star formation (SF) related feedback mechanisms represent a promising secondary ionisation source of the eDIG. We present a suite of models integrating ionisation mechanisms arising from fast shocks and photoionisation associated with star formation. When applied to the classical Baldwin-Phillips-Terlevich (BPT) diagrams, these models reveal that the ionisation budget of the eDIG ranges from 20% to 50% across our sample, with local variations of up to 20% within individual galaxy halos. This contribution correlates with the presence of filaments and other structural components observed within galaxy halos. The presence of shocks is additionally supported by the observation of high-density, high [O I] /H ratios, characteristic of shock-compressed ionised gas, likely induced by feedback from regions of intense SF within the galactic disk. These results demonstrate consistency across all galaxies analysed in this sample.
@@
To understand the physical mechanisms involved in feedback processes and the baryon cycle, which are fundamental to galaxy evolution it is essential to study the different gas phases present in the interstellar medium (ISM, Ceverino & Klypin, 2008, 2009; Hopkins et al., 2014; Klessen & Glover, 2016; Gatto et al., 2017; Grisdale, 2017; Barnes et al., 2023). One of these gas phases is known as warm ionised medium (WIM, in the context of the Milky Way) or diffuse ionised gas (DIG, for extragalactic sources; Hoyle & Ellis 1963; Hewish et al. 1968; Reynolds 1971; Reynolds et al. 1973; Guélin 1974; Reynolds 1985, 1989; Kennicutt et al. 1989; Ferguson et al. 1996). In the Milky May, this gas was found to be warm ( K) and with low density (cm-3), representing one of the most prominent components of the ISM, constituting the 20% of the ISM in volume and the 90% of the ionising hydrogen in mass (Greenawalt et al., 1997; Wang et al., 1997; Zurita et al., 2000; Otte et al., 2001; Finkbeiner, 2003; Hoopes & Walterbos, 2003; Haffner et al., 2009a). However, the DIG is typically defined by its optical emission characteristics rather than its thermodynamics (e.g. see Haffner et al. 2009b or Dirks et al. 2023). In particular, the DIG component in extragalactic halos, also known as extraplanar diffuse ionised gas (eDIG), plays a crucial role in the physical processes driving the exchange of gas, metals, and energy between the galactic disk and halo, influencing the evolution of disk galaxies (Tenorio-Tagle & Bodenheimer, 1988; Rand et al., 1990; Dettmar, 1990; Miller & Veilleux, 2003). Studies of the eDIG through imaging encompass explorations using different wavelengths (Sardaneta et al., 2024). In the optical range, the first detections of a diffuse component in the halo were made using H images (Dettmar, 1990; Rand et al., 1990; Pildis et al., 1994; Rand, 1997, 2000; Tacchella et al., 2022), These studies found that in a wide sample of star-forming galaxies the eDIG H emission in the halo correlates with the star formation rate (Rossa & Dettmar, 2003a, b; Jo et al., 2018; Lu et al., 2023). Additionally, sub-arcsecond resolution H imaging studies of individual galaxies (Howk & Savage, 2000; Rossa et al., 2004, 2005) reveal that the eDIG exhibit complex morphological structures such as filaments, superbubbles or chimneys (also detected in X-rays and radio-emission; Tüllmann et al. 2000, 2006). Alternatively, long-slit spectroscopy studies (in combination with photometric data) have enabled kinematic analyses and investigations into the potential ionisation mechanisms of the eDIG. The behaviour of the main ionisation species reveals that the [N II] /H , [S II] /H , [O III] /H and [O I] /H ratios tends to increase as a function of the distance from the midplane (Rand et al., 1990; Dettmar, 1998; Rand, 1998, 2000; Collins & Rand, 2001). Additionally, this higher degree of ionisation of the collisional lines usually implies higher electron temperatures in the eDIG compared to the H II regions (Haffner et al., 1999, 2009b; Boettcher et al., 2019). To explain these features and the increase in the emission of high ionisation species, several additional heating sources have been proposed (Haffner et al. 2009b and references therein). To infer the possible physical processes and the roles played by different morphological structures in the halo leading to the eDIG, the simultaneous spatial and spectral information provided by integral field spectrographs (IFS) has resulted in significant progress in this area. One of these studies, carried by the Sydney-AAO Multi-object Integral field spectrograph (SAMI; Croom et al. 2012; Bryant et al. 2015), used a sample of 40 low resolution edge-on galaxies (0.8 kpc ¡ 2.8 kpc) to perform a systematical exploration of the eDIG (Ho et al., 2016) noticing that the disc-halo interaction can influence the eDIG emission through starburst-driven winds. On the other hand, studies with the Mapping Nearby Galaxies at Apache Point (MaNGA; Bundy et al. 2015) IFS using also low resolution large samples found an increment of the [N II] /H indicating a temperature gradient in the halo, as well as a flattening in the H surface brightness ( ) comparable with the [O II] emission at higher distances (6 kpc).
@@
The most recent studies with the Calar Alto Legacy Integral Field Area Survey (CALIFA; Sánchez et al. 2012, 2016) studied the eDIG kinematically using a sample of 25 galaxies with spatial resolution of 0.8 kpc on average (Levy et al., 2019), finding a ”lag” in the rotation velocity of the ionised gas, i.e., a decrement of the radial velocity with the increasing of the distance from the midplane (also reported by other authors, e.g., Rand 2000; Miller & Veilleux 2003; Bizyaev et al. 2017). Nevertheless, they did not find correlations between the radial variation of the lag and the origin of the extraplanar gas. However, performing diagnoses of the main line ratios, they identified the primary ionisation sources of the eDIG the leaking of Lyman continuum (Lyc) photons from the OB stars of the H II regions, with a low contribution of an additional source in the form of hot low mass evolved stars (HOLMES; Flores-Fajardo et al. 2011; Rautio et al. 2022).
@@
In this paper we present the second study of the Bidimensional Exploration of the warm-Temperature Ionised gaS (BETIS) project, focused in the characterisation of the eDIG and the exploration of its possible ionisation mechanisms for a sample of edge-on galaxies. The paper is structured as follows. Section 2 presents and characterises the sample of edge-on galaxies, as well as describes the procedure followed for adaptive binning and the extraction of emission line features, based on the methodology developed in the first part of the series (González-Díaz et al. 2024, BETIS I henceforth). Section 3 shows a description of the ionisation structure and features found in the halo of IC1553 and the connection of the eDIG with the disc for all the sample. Section 4 presents a set of star formation and fast shocks hybrid models to explain the ionisation structure found the halo. In section 5 we show the results of applying the methodology and models of section 4 for the rest of the sample and finally, section 6 outlines the main findings in this paper.
@@
| Galaxy | RA (J2000) | DEC (J2000) | Type | z | D | P.A. | FWHM | Incl. | log SB lim. | Bin size | LeDIG(H ) | log SFRDisc |
| (deg) | (deg) | (Mpc) | (°) | (pc) | (°) | (erg/s/kpc2) | (pc) | (erg/s) | (M⊙/yr) | |||
| IC217 | 34.0435 | -11.9267 | Scd | 0.00630 | 24.32 | 35.7 | 117.91 | 82.6 | 35.6 | 94.3 | 39.11 | 0.063 |
| PGC28308 | 147.5582 | -12.0576 | Scd | 0.00907 | 45.22 | 125.2 | 153.46 | 85.5 | 35.2 | 175.4 | 39.12 | 0.084 |
| PGC30591 | 156.3603 | -15.3492 | Sd | 0.00676 | 35.22 | 169.2 | 187.83 | 86.6 | 36.0 | 136.6 | 39.13 | 0.055 |
| ESO544-27 | 33.2277 | -19.3168 | Sb | 0.00818 | 32.81 | 153.3 | 159.07 | 90.0 | 35.0 | 127.2 | 38.92 | 0.024 |
| IC1553 | 8.1671 | -25.6075 | Irr | 0.00979 | 32.25 | 15.0 | 171.99 | 78.6 | 36.5 | 125.1 | 39.80 | 0.229 |
| ESO443-21 | 194.9412 | -29.6002 | Scd | 0.00941 | 40.31 | 160.8 | 195.45 | 79.0 | 36.6 | 156.4 | 39.71 | 0.184 |
| ESO469-15 | 347.2317 | -30.8579 | Sb | 0.00545 | 19.81 | 149.2 | 86.44 | 90.0 | 35.9 | 76.9 | 38.90 | 0.029 |
| ESO157-49 | 69.9036 | -53.0126 | Sc | 0.00559 | 24.82 | 30.4 | 96.26 | 79.3 | 35.3 | 96.3 | 39.46 | 0.060 |
@@ 2 The data
For this reason, as in BETIS I, we use MUSE IFS data to undertake this task. In particular, the sample of Com19 aligns well with the previous criteria. The sample, listed in table 1, consists on eight edge-on (i ¿ 75°) nearby galaxies (z ¡ 0.01, D ¡ 46 Mpc) with high spatial resolution (full width at half maximum (FWHM) ¡ 200 pc) from the 096.B-0054(A) and 097.B-0041(A) very-large telescope (VLT) programmes. Table 1 shows the integrated star formation rate (SFR) in the disc spans between 0.024 M⊙/yr ¡ SFR ¡ 0.229 M⊙/yr. It was obtained using the Kennicutt & Evans (2012) relation: log(SFR) = log() - 41.27, with as the H luminosity, corrected for interstellar extinction assuming the Cardelli extinction law, assuming RV = 3.1 (Cardelli et al., 1989) and a Balmer decrement H /H = 2.87. The galaxies are then normal-disc galaxies with low star formation. This is consistent with the classification of these galaxies in the star formation main sequence, with specific SFR ranging between -10.5 ¡ log(sSFR/yr)¡ -9.92, corresponding barely to star-forming galaxies (using MUSE data and 3.6 m stellar masses; Muñoz-Mateos et al. 2015; Zaragoza-Cardiel et al. 2020; Rautio et al. 2022).
@@
This dataset has previously been used to analyse the eDIG in prior studies. Rautio et al. (2022) employed five of these galaxies, in conjunction with H and GALEX-FUV imaging, while Dirks et al. (2023) exclusively utilised IC1553 for a more comprehensive examination of the eDIG ionisation mechanisms within this galaxy. In our paper, we use all eight galaxies from this dataset222We downloaded the reduced datacubes directly from the ESO archive science portal. (eBETIS sample henceforth) to perform an exhaustive spectral and spatial exploration of the eDIG structure and ionisation for every galaxy in particular.
To conduct this exploration, we first extract the emission line features of the eBETIS sample in the form of binned emission line maps, following the same procedure described in BETIS I: we employ our own adaptive binning algorithm to the original datacubes, setting a S/N target of 10 based on [S II] , since this line is one of its main features of the eDIG and DIG in general, as well as it typically exhibits low S/N in both planar DIG and eDIG. Subsequently, we perform a simple stellar population (SSP) synthesis to the binned datacube using STARLIGHT (Cid Fernandes et al., 2005) to the integrated spectra of each bin. This is followed by Gaussian fitting to every emission line of interest present in each resulting nebular spectrum (see BETIS I for more details). This procedure finally results in the binned emission line maps for the H , H , [O III] , [N II] , [S II] and [O I] lines. The average sensitivity of the sample is 35.7 log(erg/s/kpc2), obtained as the 2 of the distribution of H surface brightness ( ) of the bins of every H map. The average physical size of the bins is below 175 pc. Since the average point-spread function (PSF) FWHM of the sample is below 200 pc, the binning does not significantly restrict the resolution in our galaxies.
Figure 1 shows the binned H maps of the eBETIS sample. For aesthetic purposes and to highlight the contrast between the galactic plane and the complex structure of the extraplanar gas, the plane of the galaxies is shown in greyscale. To define the galactic plane, we construct an elliptical mask by conducting a 2D Gaussian fitting to the binned H image using the python routine scypy.opt.curve_fit 333scypy.opt.curve_fit documentation., and centring the ellipse in the coordinates of the centre of the galaxy, which also coincides with the centre of the rotation curve of the galaxy. The midplane is then set at z = 0, being z the distance from the midplane, that corresponds to the centre of the minor axis of the ellipse fitted. The integrated H luminosity of the eDIG for all the galaxies is listed in Table 1. It was obtained integrating the H flux outside the ellipses previously described, and we obtain an average value of 2.51039 erg/s, that is consistent with values obtained by previous authors (Rand, 2000; Miller & Veilleux, 2003; Jo et al., 2018).